# Deep Generative Models for Detector Signature Simulation: An Analytical Taxonomy

### v1

@@ 3.1 Deep Generative Models
Despite the interoperability of Diffusion Models and their remarkable results in natural images domain [53, 54, 55], they are unable to learn a representation and like flow-based models, there is no bottleneck on the way which makes them slow-samplers. Recently, Diffusion-based models have been adapted to approximate priors in VAEs [56, 57]. A downside to consider when applying diffusion models to high-granularity detector signatures is the computational cost. The diffusion process involves multiple steps, each of which typically requires its own round of computation, making them somewhat slow samplers and computationally intensive.
@@
DGMs
Training
Likelihood Estimation
Granularity
Sampling Time
Compression
Representation Learning
VAEs
Stable
Approximate
High
Fast
Lossy
Flows
Stable
Exact
Low
Slow/Fast
Lossless
GANs
Unstable
Implicit
Ultra-High
Fast
Lossy
ARMs
Stable
Exact
Low/Mid
Slow
Lossless
Diffusion
Stable
Approximate
High
Slow
Lossy
@@ 3.1.1 Sampling Strategies
After learning a latent space (or initial inputs in the case of ARMs) for representing the input data, the generative model samples from the learned (or latent) distribution during the inference (i.e. sampling). The sampling strategies could be divided into two categories: random sampling and controllable sampling. Random sampling simply draws latent samples from a simple prior distribution, in which the model learns to approximate the distribution of the observed detector responses. The latter, on the contrary, samples detector signatures with controls over the desired properties (e.g., detector geometries, kinematic parameters, energies, luminosities, and beam parameters). Therefore, random sampling is relatively trivial, while controllable sampling usually requires extra effort in algorithm design.
Controllable generation usually manipulates the randomly sampled $z\sim p(z)$ or the encoded vector $z\sim p(z,\mathbf{L},\mathbf{C})$ in the latent space to obtain a final representation vector $z$, which is later decoded to the detector response representation $\mathbf{H}$ with expected properties. There are three types of commonly used approaches:
  • Latent Disentanglement Methods: Disentangled latent methods factorize the latent vector $z$ to $n$ parts, each of which focuses on one property $p_{n}$, following the disentanglement regularization methods that encourages the learned latent variables to be disentangled from each other. Therefore, varying one latent dimension $z_{n}$ of the latent vector $z$ will lead to property change in the generated detector responses.
  • Control Code Methods: This method incorporates a conditional vector $c$ (discrete or continuous) that explicitly controls the property of generated detector elements. In this case, the final latent representation is usually a modulation of $z$, $\mathbf{L}$ and $\mathbf{C}$. This method is the most common conditional sampling in detector signature simulation due to its versatility.
  • Iterative Editing Methods: Edit-based methods focus on iterative refinements and involve a multi-step process to modify the embedding of latent vector $z$ with conditional priors $c$ to achieve the desired characteristics in the generated detector response. Unlike disentangled sampling or conditional sampling, where the modification is done in a single pass, iterative editing typically uses optimization-based strategies (e.g., Slot-Attention mechanism [58, 59]) to iteratively update the embedding of the latent vector. This enables fine-grained control of the generated data.
@@ 3.1.2 Generation Strategies
The decoder restores the latent vector back to the data manifold. Due to the sparse, often high dimensional, and unordered nature of detector signatures, the resulting outcome of the decoder faces challenges in accurately reconstructing the data manifold or generating new signatures. This leads to artifacts or inaccuracies in the generated detector responses. Thus, existing works take three types of generation strategies for detector response generation: one-shot generation, sequential generation, and zero-shot generation.
@@
One-shot generation. One-shot generation usually generates an event in one single step. It is achieved by feeding the latent representations to neural networks to obtain the desired representation. In practice, various neural networks could be utilized, including 2D Locally Connected Networks [5], 2D and 3D Convolutional Neural Networks (CNN), Graph Neural Networks (GNN), Multi-layer Perceptron (MLP), or any combination of these modules according to different types of detectors and representations to be generated. The advantage of one-shot generation is that it generates the whole event data in a single step.
@@
If the detector topology has a non-sequential ordering (non-cylindrical/cubical topology) or irregular geometry (where sensor sizes vary with the detector depth as well as within a single layer), it is not always feasible to treat the sensor information, $\mathbf{L}$ the same as the hit manifold $\mathbf{H}$. Consequently, using a grid-based (homogenous) representation to characterize the full detector signatures is not the most efficient method. Instead, a variable multi-set (heterogeneous) representation is preferable. Detector signatures in multi-set representation can be related similarly to the Mesh representation of 3D objects, which consists of a collection of vertices (like detector hits) and polygon faces (like layer information). For example, the Pixel Vertex Detector (PXD) [35] detector at Belle II carries such a non-trivial topology where the detector is a Toroid with an octagon inner layer and a dodecagon outer layer. Or the ATLAS experiment [60] electromagnetic calorimeter incorporates an irregular geometry of calorimeter segmentation. As a result, in the one-shot generation of multi-set representations of detector elements, it is still of paramount importance to incorporate suitable inductive biases to capture the detector geometry and topology.
@@
Sequential generation. In contrast to one-shot generation, sequential generation generates the detector responses consecutively in a few steps. As there is ordering naturally defined for detector layers such as a calorimeter detector, sequential generation has to follow a certain sequential inductive bias for the generation. This is usually done by generating probabilistic sensor or layer features while sampling and feeding step-by-step the reconstructed detector response following a predefined ordering. Despite slow sampling, sequential generative models enjoy the benefit of auto-regressive and causal reasoning, which prompts a precise correlation modeling of the data. Therefore, it could be easily combined with constraints in each of the generation steps when the responses to be generated should obey certain restrictions. Using sequential sampling, when generating detector signatures with either high-granularity or long detector layers, the error will accumulate at each step, possibly resulting in discrepancies in the final generated and observed detector signatures. Also, in the case of non-cylindrical detector topologies such as PXD at Belle II, pure autoregressive reasoning could introduce a non-existing sequential bias into the generated data. In this case, one has to change the perspective from the single “shower” generation to the full “event” generation where correlations between different directions (various angles) become important.
@@
Zero-shot generation. Zero-shot detector response generation refers to the process of generating detector responses without any prior exposure to the specific representation or structure of the “new” target detector signatures. The new target detector signatures could belong to beam parameters beyond the training range, higher incident energies, higher luminosities, or detector responses for uninstalled sensors, as will be discussed in Section 4.3. The term Zero-Shot comes from the concept of zero-shot learning, which is when a model can recognize or generate outputs for new, unseen categories or tasks without any training examples. This is achieved by leveraging the model’s pre-existing knowledge and generalization abilities, typically acquired during the pre-training phase on large datasets by incorporating symmetries and constraints directly into the generative model. This strategy leverages the latent space’s ability to capture the essential characteristics of the data manifold, enabling the generation of plausible detector responses for regions and conditions beyond the current data at hand. The core idea that is introduced in this study is to extrapolate to the Out-Of-Distribution (OOD) data by designing a model to be more flexible and adaptive, allowing it to accommodate detector geometries, representations, and conditions beyond the training data.
3.1.3 Representative Work and Datasets
@@
In this section, we discuss representative works for deep generative models in detector signature simulation with an emphasis on how they handle generation and sampling. These works deal with a wide variety of representations for different detector signatures, including Shower Simulation, Jet Simulation, and track (Hit) Simulation — that yield datasets with different granularity and complexity.
@@
Shower signatures correspond to datasets that encapsulate the complex dynamics of particle showers in calorimeters. These datasets often include the spatial distribution of energy deposits, the type of interacting particle, incident angle, and other features that distinguish electromagnetic from hadronic showers. Depending on the calorimeter design and the simulation framework, the granularity can range from very fine to coarser bins that encapsulate multiple layers of the calorimeter. The highest granularity studied by the surrogate models for calorimeter showers was by Khattak et al. [61, 62] with 65k information channels.
@@
Jets are sprays of particles that are produced when high-energy quarks or gluons (partons) propagate through the detector. These partons undergo a parton shower and hadronization process, which results in a jet of hadrons. Jet signatures contain not only the final state particles in a jet but also their kinematic properties and relationships, enabling a multidimensional view of jets. The granularity in jet signatures is usually more variable, often depending on the event’s complexity. Kinematic properties like transverse momentum, azimuthal angle, and pseudo rapidity can be considered at various resolutions, depending on the level of detail required for the specific analysis.
@@
Track signatures come from the direct detector hits or fired pixels. They contain the signals corresponding to charged particle interactions with the detector elements. They often have high granularity, mimicking the detector’s actual resolution, which can be on the scale $\mathcal{O}(10^{7})$ channels. Each hit usually contains information like position, time, and the deposited charge, offering a granular view suitable for the downstream track reconstruction algorithms. The highest granularity studied by the surrogate models for charged tracks was by Hashemi et al. [63, 45] with 7,680,000 information channels per event.
@@
In general, based on granularity, one can categorize DGMs into $4$ classes: low granularity data with $\mathcal{O}(100)$ channels like jets, mid granularity data with $\mathcal{O}(10^{4})$, high granularity data with $\mathcal{O}(10^{5})$ channels, and ultra-high granularity data with $\mathcal{O}(10^{7})$ input channels such as High Granularity Calorimeter (HGCAL) [33] at CMS [64] with roughly $6$ million channels, the PXD at Belle II [36] with more than $7.5$ million-pixel channels, or the EPICAL-2 [65] ultra-high granularity electromagnetic calorimeter with $12.5$ million-pixel channels. In Table 2, we provide the top $15$ SOTA generative models with the highest granularities ever studied.
@@
Historically, each newly introduced architecture was focused on a specific detector layout and hence worked with its own dataset. Such a situation is unfavorable for a large-scale and inter-experiment deployment of DGMs for detector simulation, as it is almost impossible to study improvements of new architectures over older ones. To alleviate this problem, a public data challenge called the Fast Calorimeter Simulation Challenge 2022 — CaloChallenge [66] was created. It consists of simulated showers of increasing dimensionality, ranging from $\mathcal{O}(10^{2})$ to $\mathcal{O}(10^{5})$ dimensions. These datasets can serve as a benchmark for the existing and future DGMs for mid to high-granularity experiments and will help to improve our understanding of common struggles, advantages, and disadvantages, as well as the scaling behavior of different DGM approaches.
@@
In addition, there are other public datasets with different granularities coming from various experiments. The Electromagnetic Calorimeter Shower dataset [67, 68] was the first public dataset on calorimeter showers corresponding to energy deposits from positrons, photons, and charged Pions. Diefenbacher et al. [69, 70] also provided a public dataset for high-resolution photon showers conditioned on the incident energy and angle. JetNet [71] is another important benchmarking contribution to jet datasets, represented as multi-sets containing gluon, top quark, and light quark jets. JetClass [72, 73], a more complex public jet dataset, contains 10 classes of jets carrying three categories of features, kinematics, particle identification, and trajectory displacement information. Hashemi et al. [74, 45] also recently introduced a public dataset for ultra-high granularity detector signatures with a non-trivial detector topology of PXD at Belle II, conditioned on the sensor position. (radius and angle). This dataset acts as a benchmark for evaluating the DGMs geared towards real experiments with ultra-high granularity and provides insights into the scalability of current DGM approaches towards the HL-LHC era. To spur further research and development, frameworks such as COCOA [75] and the Open Data Detector [76, 77] have been introduced recently.

### v2

@@ 3.1 Deep Generative Models
Along with the interoperability of diffusion models and their remarkable results in natural images domain [ho_denoising_2020, kingma_variational_2023, ramesh_hierarchical_2022], recent studies [2024arXiv240700783F] are showing the advancement of representation learning with them unlike flow-based models where there is no information bottleneck (the latent space has the same dimension as the data space). Moreover, diffusion-based models can be adapted to approximate priors in VAEs [vahdat_score-based_2021, wehenkel_diffusion_2021]. A downside to consider when applying diffusion models to high-granularity detector signatures is the computational cost. The diffusion process involves multiple steps, each of which typically requires its own round of computation, making them somewhat slow samplers and computationally intensive.
@@
DGMs Training Likelihood Estimation Granularity Sampling Time Compression Representation Learning
VAEs Stable Approximate High Fast Lossy
Flows Stable Exact Low Slow/Fast Lossless
GANs Unstable Implicit Ultra-High Fast Lossy
ARMs Stable Exact Low/Mid Slow Lossless
Diffusion Stable Approximate High Slow Lossy
@@ 3.1.1 Sampling Strategies
After learning a latent space (or initial inputs in the case of ARMs) for representing the input data, the generative model samples from the learned (or latent) distribution during the inference (i.e. sampling). The sampling strategies could be divided into two categories: random sampling and controllable sampling. Random sampling simply draws latent samples from a simple prior distribution (Normal Distribution), in which the model learns to approximate the distribution of the observed detector responses. This corresponds to sampling from the distribution $p(\mathbf{D})=p(\mathbf{L},\mathbf{C},\mathbf{H})$. The latter, along with random noise sampling, samples detector signatures with controls over the desired properties (e.g., detector geometries, kinematic parameters, energies, luminosities, and beam parameters) as well. This corresponds to sampling from the conditional distribution $p(\mathbf{H}|\mathbf{L},\mathbf{C})$ or $p(\mathbf{H},\mathbf{L}|\mathbf{C})$, depending on the situation. Therefore, random sampling is relatively trivial, while controllable sampling usually requires extra effort in algorithm design.
Controllable sampling usually manipulates the randomly sampled $z\sim p(z)$ or the encoded vector $z\sim p(z,\mathbf{L},\mathbf{C})$ in the latent space to obtain a final representation vector $z$, which is later decoded to the detector response representation $\mathbf{H}$ with expected properties. There are three types of commonly used approaches:
  • Latent Disentanglement Methods: Disentangled latent methods [bengio_representation_2014, kim_disentangling_2019, mathieu_disentangling_2019, collins_exploration_2022] factorize the latent vector $z$ to $n$ parts, each of which focuses on one property $p_{n}$, encouraging the learned latent variables to be disentangled from each other. Therefore, varying one latent dimension $z_{n}$ of the latent vector $z$ will lead to property change in the generated detector responses.
  • Control Code Methods: This method [sohn_learning_2015, mirza_conditional_2014, winkler_learning_2023] incorporates a conditional vector $c$ (discrete or continuous) that explicitly controls the property of generated detector elements. In this case, the final latent representation is usually a modulation of $z$, $\mathbf{L}$ and $\mathbf{C}$. This method is the most common conditional sampling in detector signature simulation due to its versatility.
  • Iterative Editing Methods: Edit-based methods [hjelm_iterative_2018, grover_graphite_2019, cheng_controllable_2020, winterhalder_latent_2021] focus on iterative refinements and involve a multi-step process to modify the embedding of latent vector $z$ with conditional priors $c$ to achieve the desired characteristics in the generated detector response. Unlike disentangled sampling or control code methods, where the modification is done in a single pass, iterative editing typically uses optimization-based strategies (e.g., Slot-Attention mechanism [locatello_object-centric_2020, di_bello_conditional_2022]) to iteratively update the embedding of the latent vector. This enables fine-grained control of the generated data.
@@ 3.1.2 Generation Strategies
The generator (in some models this is done by the decoder) restores the latent vector back to the data manifold. Due to the sparse, often high dimensional, and unordered nature of detector signatures (before fixing the representation with voxelization or pixelization), the resulting outcome of the generator / decoder faces challenges in accurately reconstructing the data or generating new signatures. This could lead to artifacts or inaccuracies in the generated detector responses, if not handled carefully. Thus, existing works take three types of generation strategies for detector response generation: one-shot generation, sequential generation, and zero-shot generation.
@@
One-shot generation. In one-shot generation [rezende_one-shot_2016, kipf_variational_2016, wang_graphgan_2017, de_cao_molgan_2022], an event is generated in one single step. It is achieved by feeding the latent representations to neural networks to obtain the desired representation. In practice, various neural networks could be utilized, including two-dimensional (2D) Locally Connected Networks [de_oliveira_learning_2017], 2D and three-dimensional (3D) Convolutional Neural Networks (CNN), Graph Neural Networks (GNN), Multi-layer Perceptron (MLP), or any combination of these modules according to different types of detectors and representations to be generated. The advantage of one-shot generation is that it generates the whole event data in a single step.
@@
If the detector topology has a non-sequential ordering (non-cylindrical/cubical topology) or irregular geometry (where sensor sizes vary with the detector depth as well as within a single layer), it is not always feasible to treat the sensor information, $\mathbf{L}$ the same as the hit manifold $\mathbf{H}$. Consequently, using a grid-based (homogeneous) representation to characterize the full detector signatures is not the most efficient method. Instead, a variable multi-set (heterogeneous) representation is preferable. Detector signatures in multi-set representation can be related similarly to the Mesh representation of 3D objects, which consists of a collection of vertices (like detector hits) and polygon faces (like layer information). For example, the Pixel Vertex Detector (PXD) [giakoustidis_status_2023] detector at Belle II carries such a non-trivial topology where the detector is a Toroid with an octagon inner layer and a dodecagon outer layer. Or the ATLAS experiment [noauthor_atlas_nodate] electromagnetic calorimeter incorporates an irregular geometry of calorimeter segmentation. As a result, in the one-shot generation of multi-set representations of detector elements, it is still of paramount importance to incorporate suitable inductive biases to capture the detector geometry and topology.
@@
Sequential generation. In contrast to one-shot generation, sequential generation [bacciu_edge-based_2020, liao_efcient_nodate, ingraham_generative_nodate] generates the detector responses consecutively in a few steps. As there is ordering naturally defined for detector layers such as a calorimeter detector, sequential generation has to follow a certain sequential inductive bias for the generation. This is usually done by generating probabilistic sensor or layer features while sampling and feeding step-by-step the reconstructed detector response following a predefined ordering. Despite slow sampling, sequential generative models enjoy the benefit of auto-regressive and causal reasoning, which prompts a precise correlation modeling of the data. Therefore, it could be easily combined with constraints in each of the generation steps when the responses to be generated should obey certain restrictions. Using sequential sampling, when generating detector signatures with either high-granularity or long detector layers, the error will accumulate at each step, possibly resulting in discrepancies in the final generated and observed detector signatures. Also, in the case of non-cylindrical detector topologies such as PXD at Belle II, pure autoregressive reasoning could introduce a non-existing sequential bias into the generated data. In this case, one has to change the perspective from the single “shower” generation to the full “event” generation, where correlations between different directions (various angles) become important. Note that while ARMs rely on sequential generation by construction, the sequential generation strategy is more general an can also be applied to other DGM architectures, for example to break down the problem into easier tasks.
@@
Zero-shot generation. Zero-shot detector response generation refers to the process of generating detector responses without any prior exposure to the specific representation or structure of the “new” target detector signatures [hashemi_deep_2024]. The new target detector signatures could belong to beam parameters beyond the training range, higher incident energies, higher luminosities, or detector responses for uninstalled sensors, as will be discussed in Section 4.3. The term Zero-Shot comes from the concept of zero-shot learning [noauthor_deep_nodate], which is when a model can recognize or generate outputs for new, unseen categories or tasks without any training examples. This is achieved by leveraging the model’s pre-existing knowledge and generalization abilities, typically acquired during the pre-training phase on large datasets by incorporating symmetries and constraints directly into the generative model. This strategy [liu_attribute_2020, li_graph_2023, chan_deep_2021, padmakumar_extrapolative_2023] leverages the latent space’s ability to capture the essential characteristics of the data manifold, enabling the generation of plausible detector responses for regions and conditions beyond the current data at hand. The core idea that is introduced in this study is to extrapolate to the Out-Of-Distribution (OOD) data by designing a model to be more flexible and adaptive, allowing it to accommodate detector geometries, representations, and conditions beyond the training data.
3.1.3 Evaluation
The evaluation of DGMs is an active area of research in which several quantitative and qualitative measures have been proposed so far [krause_caloflow_2023-1, hashemi_ultra-high-resolution_2023, kansal_evaluating_2023, acosta_comparison_2023, hashemi_deep_2024, das_how_2023]. Such an evaluation can be done at different levels, each with it’s own advantages and disadvantages. In addition to metrics assessing the quality of generated samples, one should also evaluate the time required to generate the samples, as there is usually a trade-off between these two aspects [xiao_tackling_2022]. While a detailed comparison of all DGMs presented here is not possible, a subset of them will be compared to each other as part of the CaloChallenge [michele_fast_2022] in a dedicated summary publication [CaloChallenge_writeup].
Histogram or Marginal Distribution-based Methods (signature-level). The comparison of samples (or quantities derived from them) between the DGM and the reference dataset in histograms is a fast way to evaluate the quality of a DGM. However, this method only provides a necessary, not a sufficient condition, as it is just a one-dimensional projection of a higher-dimensional space and therefore blind to correlations between the features.
Downstream Physics-based Methods. From the physics perspective, we are interested in the difference between using samples from the DGMs vs. using samples from the traditional simulation chain. Therefore, instead of comparing the full distributions, one only focuses on the derived observables one actually cares about. While this method of evaluation the most aligned to our physics goals, it comes at the drawback that the selection of observables is fixed at the point of evaluation. The inclusion of additional observables at a later stage might reveal a suboptimal DGM. Some works included such a methods in their studies, investigating various reconstructed parameters such as the tracking impact parameters [hashemi_ultra-high-resolution_2023] or shower reconstructed energies [buhmann_hadrons_2021].
Neural network (NN)-based Methods. These methods use the power of neural networks to learn correlations in the high-dimensional data, offering methods of evaluation that are sensitive to the entire distribution.
One way of using NN-based methods is to do the evaluation of DGMs as two sample test. One can use the Neyman-Pearson lemma [noauthor_ix_nodate] and train a neural classifier to distinguish samples from the DGM from samples of the reference data [lopez-paz_revisiting_2018]. If a powerful classifier is not able to distinguish the two sets, one concludes that the two underlying distributions are identical. This approach provides not only a single number (Area Under the receiver operating characteristic Curve (AUC)) indicating the quality of the DGM, but also offers a way of analyzing which regions of phase space have problems (via the distribution of the classifier outputs [das_how_2023]). The classifier-based test [krause_caloflow_2023-1] gives the most powerful result on the two-sample test. However, while these results are sufficient for the comparison, they are not always necessary. For example, the disagreement between the DGM and the reference distribution could be in an irrelevant region of phase space.
Another way of incorporating the NN-based methods is the feature space-matching methods [borji_pros_2022] that capture more complex aspects of the data distributions, including the diversity, patterns, and higher-level attributes of the data. They rely on a pre-trained (on the real data) multi-class classifier (on the conditions $C$) with any backbone model [salimans_improved_2016, zietlow_demystifying_2021, chong_effectively_2020, parmar_aliased_2022]. Then, one can calculate a distance function between the embedding space (feature space) of the backbone network of the generated and the real data. The lower this distance, the higher the fidelity of the generated data [betzalel_study_2022]. In detector signature simulation, these metrics have shown a very high sensitivity to small changes in the real data distribution’s quality [hashemi_ultra-high-resolution_2023, kansal_evaluating_2023] and diversity [hashemi_deep_2024].
3.2 Representative Work and Datasets
@@
In this section, we discuss representative works for deep generative models in detector signature simulation with an emphasis on how they handle generation and sampling. For each model, we first go through the works that target detector response (shower and track) generation, then we review models that aim to do event or jet simulation.
@@
Shower signatures correspond to datasets that encapsulate the complex dynamics of particle showers in calorimeters. These datasets often include the spatial distribution of energy deposits, the type of interacting particle, incident angle, and other features that distinguish electromagnetic from hadronic showers. Depending on the calorimeter design and the simulation framework, the granularity can range from very fine to coarser bins that encapsulate multiple layers of the calorimeter.
@@
Jets are collimated sprays of particles that are produced when high-energy quarks or gluons (partons) undergo a hadronization process. Jet signatures contain not only the final state particles in a jet but also their kinematic properties and relationships, enabling a multidimensional view of jets. The granularity in jet signatures is usually more variable, often depending on the event’s complexity. Kinematic properties like transverse momentum, azimuthal angle, and pseudo rapidity can be considered at various resolutions, depending on the level of detail required for the specific analysis.
@@
Track signatures come from the direct detector hits or fired pixels. They contain the signals corresponding to charged particle interactions with the detector elements. They often have extremely high granularities, mimicking the detector’s actual resolution, which can be on the scale of more than $\mathcal{O}(10^{8})$ channels. Each hit usually contains information like position, time, and the deposited charge of the hit, offering a granular view suitable for the downstream track reconstruction algorithms.
@@
In general, based on granularity, one can categorize DGMs into $4$ classes: low granularity data with $\mathcal{O}(100)$ channels like jets, mid granularity data with $\mathcal{O}(10^{4})$ channels, high granularity data with $\mathcal{O}(10^{5})$ channels, and ultra-high granularity data with up to $\mathcal{O}(10^{7})$ input channels such as High Granularity Calorimeter (HGCAL) [magnan_hgcal_2017] at CMS [noauthor_phase-2_2017] with roughly $6$ million channels, the PXD at Belle II [abe_belle_2010] with more than $7.5$ million-pixel channels, or the future EPICAL-2 [noauthor_results_nodate] ultra-high granularity electromagnetic calorimeter with $12.5$ million-pixel channels at ALICE [aamodt_alice_2008]. In Table 2, we provide the top $15$ state-of-the-art (SOTA) generative models with the highest granularities ever studied.
@@
Historically, each newly introduced architecture was focused on a specific detector layout and hence worked with its own dataset. Such a situation is unfavorable for a large-scale and inter-experiment deployment of DGMs for detector simulation, as it is almost impossible to study improvements of new architectures over older ones. To alleviate this problem, a public data challenge called the Fast Calorimeter Simulation Challenge 2022 — CaloChallenge [michele_fast_2022] was created. It consists of simulated showers of increasing dimensionality, ranging from $\mathcal{O}(10^{2})$ to $\mathcal{O}(10^{5})$ dimensions. These datasets can serve as a benchmark for the existing and future DGMs for mid to high-granularity experiments and will help to improve our understanding of common struggles, advantages, and disadvantages, as well as the scaling behavior of different DGM approaches.
@@
In addition, there are other public datasets with different granularities coming from various experiments. The Electromagnetic Calorimeter Shower dataset (CaloGAN) [nachman_electromagnetic_2017, paganini_calogan_2018] was the first public dataset on calorimeter showers corresponding to energy deposits from positrons, photons, and charged Pions. Diefenbacher et al. [noauthor_photon_nodate, diefenbacher_new_2023] also provided a public dataset (ILD dataset) for high-resolution photon showers conditioned on the incident energy and angle. JetNet [kansal_jetnet_2023] is another important benchmarking contribution to jet datasets, represented as multi-sets containing gluon, top quark, and light quark jets. JetClass [qu_particle_2022, qu_jetclass_2022], a more complex public jet dataset, contains 10 classes of jets carrying three categories of features, kinematics, particle identification, and trajectory displacement information. Hashemi et al. [hashemi_ultra-high_nodate, hashemi_ultra-high-resolution_2023] also recently introduced a public dataset (PXD dataset) for ultra-high granularity detector signatures with a non-trivial detector topology of PXD at Belle II, conditioned on the sensor position, i.e. radius and angle. This dataset acts as a benchmark for evaluating the DGMs geared towards real experiments with ultra-high granularity and provides insights into the scalability of current DGM approaches towards the HL-LHC era. To spur further research and development, frameworks such as COCOA [charkin-gorbulin_configurable_2023] and the Open Data Detector [gessinger-befurt_open_2023, noauthor_acts_nodate] have been introduced recently.